# COMPASS: Robust Feature Conformal Prediction for Medical Segmentation Metrics

### v1

@@ C.1 Algorithm 1: COMPASS
15:procedure Calibration($D_{\mathrm{cal}}$, model, $h,\alpha,\dots$)
16:  Initialize scores $\mathcal{R}\leftarrow\emptyset$.
17:  for $i=1,\dots,n$ do
18:   Define representation $\hat{z}_{i}$ and direction $\Delta_{i}$.
19:   $R_{i}\leftarrow\text{Symmetric-Binary-Search}(\hat{z}_{i},\Delta_{i},y_{i})$.
20:   $\mathcal{R}\leftarrow\mathcal{R}\cup\{R_{i}\}$.
21:  end for
22:  $\hat{\beta}\leftarrow\text{Quantile}(\mathcal{R},\frac{\lceil(1-\alpha)(n+1)\rceil}{n})$.
23:  return $\hat{\beta}$
24:end procedure
25:
26:procedure Inference($x_{n+1}$, model, $h,\hat{\beta}$$\dots$)
27:  Define representation $\hat{z}_{n+1}$ and direction $\Delta_{n+1}$.
@@
29:  return $S_{\hat{\beta}}(x_{n+1})$
30:end procedure
C.2 Algorithm 2: Asymmetric Perturbations with COMPASS
We extend COMPASS to provide separate control over the lower and upper bounds of the prediction interval—the asymmetric case. We calculate two distinct non-conformity scores ($R_{\text{lo}},R_{\text{hi}}$) and two corresponding perturbation magnitudes ($\hat{\beta}_{\text{lo}},\hat{\beta}_{\text{hi}}$) based on separate miscoverage rates ($\alpha_{\text{lo}},\alpha_{\text{hi}}$).
Algorithm 2 Asymmetric COMPASS
1:Training $D_{\mathrm{tr}}$, Calibration $D_{\mathrm{cal}}=\{(x_{i},y_{i})\}_{i=1}^{n}$, Test $x_{n+1}$, $\alpha_{\mathrm{lo}},\alpha_{\mathrm{hi}}$, $L$.
@@ C.2 Algorithm 2: Asymmetric Perturbations with COMPASS
4:procedure Training($D_{\mathrm{tr}}$, $L$) $\triangleright$ Identical to Symmetric COMPASS
5:  This stage is identical to the Training procedure in Algorithm 1.
6:end procedure
7:
8:procedure Calibration($D_{\mathrm{cal}}$, model, $h,\alpha_{\mathrm{lo}},\alpha_{\mathrm{hi}},\dots$)
9:  Initialize $\mathcal{R}_{\mathrm{lo}}\leftarrow\emptyset,\mathcal{R}_{\mathrm{hi}}\leftarrow\emptyset$.
10:  for $i=1,\dots,n$ do
11:   Define representation $\hat{z}_{i}$ and direction $\Delta_{i}$.
12:   $(R_{i,\mathrm{lo}},R_{i,\mathrm{hi}})\leftarrow\text{Asymmetric-Binary-Search}(\hat{z}_{i},\Delta_{i},y_{i})$.
13:   $\mathcal{R}_{\mathrm{lo}}\leftarrow\mathcal{R}_{\mathrm{lo}}\cup\{R_{i,\mathrm{lo}}\}$, $\mathcal{R}_{\mathrm{hi}}\leftarrow\mathcal{R}_{\mathrm{hi}}\cup\{R_{i,\mathrm{hi}}\}$.
14:  end for
15:  $\hat{\beta}_{\mathrm{lo}}\leftarrow\text{Quantile}(\mathcal{R}_{\mathrm{lo}},\frac{\lceil(1-\alpha_{\mathrm{lo}})(n+1)\rceil}{n})$.
16:  $\hat{\beta}_{\mathrm{hi}}\leftarrow\text{Quantile}(\mathcal{R}_{\mathrm{hi}},\frac{\lceil(1-\alpha_{\mathrm{hi}})(n+1)\rceil}{n})$.
17:  return $\hat{\beta}_{\mathrm{lo}},\hat{\beta}_{\mathrm{hi}}$
18:end procedure
19:
20:procedure Inference($x_{n+1}$, model, $h,\hat{\beta}_{\mathrm{lo}},\hat{\beta}_{\mathrm{hi}},\dots$)
21:  Define representation $\hat{z}_{n+1}$ and direction $\Delta_{n+1}$.
22:  $S(x_{n+1})\leftarrow[h(g(\hat{z}_{n+1}-\hat{\beta}_{\mathrm{lo}}\Delta_{n+1})),h(g(\hat{z}_{n+1}+\hat{\beta}_{\mathrm{hi}}\Delta_{n+1}))]$.
23:  return $S(x_{n+1})$
24:end procedure
C.3 Score-Finding Procedures
The core computational task in the calibration stage is finding the minimal perturbation magnitude ($\beta$) that satisfies the coverage condition. We present two search methods, linear and binary, each with symmetric and asymmetric versions.
C.3.1 Linear Search
With the linear search method, we perform a straightforward, iterative approach to finding the non-conformity score. We start with a perturbation magnitude of zero and incrementally increase it by a fixed step size, $\beta_{\text{step}}$. At each step, we perform a forward pass to check if the coverage condition is met. We terminate the search as soon as the condition is satisfied, returning the current magnitude as the score. While simple and robust, this method can be computationally expensive.
Algorithm 3 Symmetric Linear Search
1:function Symmetric-Linear-Search($\hat{z},\Delta,y$)
2:  $\beta\leftarrow 0$.
3:  repeat
4:   $\beta\leftarrow\beta+\beta_{\text{step}}$.
5:  until $y\in P(\hat{z},\Delta,\beta)$
6:  return $\beta$.
7:end function
Algorithm 4 Asymmetric Linear Search
1:function Asymmetric-Linear-Search($\hat{z},\Delta,y$)
2:  $\beta_{\text{lo}}\leftarrow 0$. Repeat $\beta_{\text{lo}}\leftarrow\beta_{\text{lo}}+\beta_{\text{step}}$ until $y\geq P(\hat{z},\Delta,\beta_{\text{lo}})_{\text{lower}}$.
3:  $R_{\mathrm{lo}}\leftarrow\beta_{\text{lo}}$.
4:  $\beta_{\text{hi}}\leftarrow 0$. Repeat $\beta_{\text{hi}}\leftarrow\beta_{\text{hi}}+\beta_{\text{step}}$ until $y\leq P(\hat{z},\Delta,\beta_{\text{hi}})_{\text{upper}}$.
5:  $R_{\mathrm{hi}}\leftarrow\beta_{\text{hi}}$.
6:  return $(R_{\mathrm{lo}},R_{\mathrm{hi}})$.
7:end function
C.3.2 Binary Search
With the binary search method, we use a highly efficient divide and conquer algorithm. We first define a search interval, $0,\beta_{\text{range}}]$, and repeatedly halve this interval by testing its midpoint. If the midpoint magnitude achieves coverage, we continue the search in the lower half; otherwise, we continue in the upper half. We repeat this process for a fixed number of iterations, $k_{\text{max}}$, allowing the search to converge on the minimal required magnitude with logarithmic time complexity.
@@ C.3.2 Binary Search
5:   if $y\in P(\hat{z},\Delta,b_{\text{mid}})$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
6:   else $b_{\text{low}}\leftarrow b_{\text{mid}}$
7:   end if
8:  end for
9:  return $b_{\text{high}}$.
10:end function
Algorithm 6 Asymmetric Binary Search
1:function Asymmetric-Binary-Search($\hat{z},\Delta,y$)
2:  Upper Bound Score ($R_{\mathrm{hi}}$):
3:  $b_{\text{low}}\leftarrow 0,b_{\text{high}}\leftarrow\beta_{\text{range}}$.
4:  for $k=1,\dots,k_{\text{max}}$ do
5:   $b_{\text{mid}}\leftarrow(b_{\text{low}}+b_{\text{high}})/2$.
6:   if $y\leq P(\hat{z},\Delta,b_{\text{mid}})_{\text{upper}}$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
@@
10:  $R_{\mathrm{hi}}\leftarrow b_{\text{high}}$.
11:  Lower Bound Score ($R_{\mathrm{lo}}$):
12:  $b_{\text{low}}\leftarrow 0,b_{\text{high}}\leftarrow\beta_{\text{range}}$.
13:  for $k=1,\dots,k_{\text{max}}$ do
14:   $b_{\text{mid}}\leftarrow(b_{\text{low}}+b_{\text{high}})/2$.
15:   if $y\geq P(\hat{z},\Delta,b_{\text{mid}})_{\text{lower}}$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
16:   else $b_{\text{low}}\leftarrow b_{\text{mid}}$
17:   end if
18:  end for
19:  $R_{\mathrm{lo}}\leftarrow b_{\text{high}}$.
20:  return $(R_{\mathrm{lo}},R_{\mathrm{hi}})$.
21:end function
C.3.3 Computational Complexity of Search Methods
The efficiency of the calibration stage is dominated by the score-finding function, as it requires repeated forward passes through the model’s decoder.
  • Linear Search: The computational complexity is $\mathcal{O}(N)$, where $N=\beta_{\text{max}}/\beta_{\text{step}}$ is the number of steps in the search. In the worst case, this requires $N$ expensive forward passes for each calibration sample.
  • Binary Search: The complexity is $\mathcal{O}(\log(N))$, where the search space is effectively reduced by half in each step. For a fixed number of iterations $k_{\text{max}}$, the number of forward passes is constant and small (e.g., 30-100), regardless of the range of $\beta$. This offers a dramatic speedup over the linear search, especially when high precision is required.
@@ Appendix D Additional Tables
D.1 Empirical Coverages
Coverage
Dataset $\alpha$ COMPASS-J COMPASS-L E2E-CQR Local Output-CQR SCP FCP
H&E 0.05 0.954±0.019 0.959±0.02 0.959±0.019 0.954±0.018 0.955±0.017 0.958±0.018 0.952±0.010
0.1 0.895±0.029 0.9±0.026 0.906±0.027 0.903±0.031 0.902±0.028 0.906±0.027 0.897±0.014
0.15 0.85±0.028 0.859±0.032 0.852±0.031 0.856±0.031 0.855±0.03 0.859±0.031 0.850±0.018
Skin Lesion 0.05 0.95±0.01 0.95±0.009 0.951±0.01 0.951±0.009 0.949±0.01 0.952±0.01 0.949±0.006
0.1 0.901±0.014 0.899±0.012 0.899±0.013 0.9±0.013 0.9±0.014 0.9±0.014 0.903±0.007
0.15 0.85±0.015 0.851±0.015 0.851±0.013 0.852±0.015 0.849±0.015 0.852±0.015 0.850±0.008
Nodules 0.05 0.951±0.017 0.951±0.015 0.953±0.018 0.951±0.018 0.949±0.019 0.955±0.017 0.950±0.010
0.1 0.899±0.026 0.902±0.024 0.906±0.026 0.902±0.026 0.901±0.022 0.907±0.027 0.899±0.014
0.15 0.85±0.032 0.85±0.033 0.857±0.033 0.848±0.031 0.848±0.032 0.852±0.027 0.856±0.016
PolyP 0.05 0.952±0.021 0.951±0.021 0.954±0.021 0.951±0.02 0.95±0.021 0.953±0.02 0.952±0.013
0.1 0.899±0.032 0.899±0.03 0.907±0.026 0.907±0.027 0.903±0.029 0.905±0.029 0.901±0.017
0.15 0.85±0.031 0.852±0.032 0.857±0.033 0.853±0.035 0.852±0.034 0.858±0.036 0.854±0.02
@@ D.1 Empirical Coverages
D.2 COMPASS versus FCP
Interval Size (Mean±STD, Pixels2)
@@ D.2 COMPASS versus FCP
H&E 0.05 4635±569 4542±449 9702±1535
0.1 2512±160 2655±205 6463±1193
0.15 2364±145 2513±145 4979±647
Skin Lesion 0.05 1590±61 1633±84 14838±10
0.1 1149±41 1170±47 12445±51
0.15 920±24 932±26 8267±79
Nodules 0.05 3184±190 3293±208 16247±2
0.1 2388±130 2469±141 16143±3
0.15 1940±106 2005±138 16056±4
PolyP 0.05 5061±497 5358±563 15937±9
0.1 3940±262 4101±286 15644±10
0.15 3228±191 3414±254 15353±14
D.3 Weighted Conformal Prediction Interval Size and Coverage
@@
H&E Class COMPASS-J 2055.0±220.0 0.911±0.024
COMPASS-L 2065.0±263.0 0.903±0.024
E2E-CQR 2444.0±278.0 0.9±0.027
Local 3113.0±201.0 0.907±0.024
Output-CQR 2890.0±265.0 0.901±0.024
SCP 2371.0±290.0 0.9±0.026
Feature COMPASS-J 1916.0±177.0 0.909±0.024
COMPASS-L 1944.0±244.0 0.896±0.024
E2E-CQR 2197.0±288.0 0.884±0.029
Jacobian COMPASS-J 1822.0±200.0 0.896±0.031
Unweighted COMPASS-J 1810.0±127.0 0.886±0.025
COMPASS-L 1759.0±159.0 0.884±0.024
E2E-CQR 1993.0±196.0 0.871±0.026
Local 2833.0±158.0 0.88±0.027
Output-CQR 2563.0±134.0 0.879±0.024
SCP 2003.0±169.0 0.878±0.024
Skin Lesion Class COMPASS-J 1318.0±132.0 0.896±0.019
COMPASS-L 1373.0±136.0 0.895±0.019
E2E-CQR 1335.0±277.0 0.887±0.021
Local 2426.0±354.0 0.893±0.02
Output-CQR 3380.0±43.0 0.888±0.018
SCP 1905.0±475.0 0.888±0.024
Feature COMPASS-J 1573.0±158.0 0.919±0.014
COMPASS-L 1669.0±168.0 0.919±0.013
E2E-CQR 2030.0±269.0 0.919±0.012
Jacobian COMPASS-J 1485.0±126.0 0.914±0.013
Unweighted COMPASS-J 2183.0±132.0 0.94±0.005
COMPASS-L 2211.0±150.0 0.939±0.007
E2E-CQR 3400.0±226.0 0.945±0.006
Local 5446.0±380.0 0.947±0.006
Output-CQR 4201.0±145.0 0.945±0.008
SCP 4707.0±323.0 0.948±0.006
D.4 Other architectures Interval Size and Coverage
Interval Size (Mean±STD, Pixels2)
@@ D.4 Other architectures Interval Size and Coverage
0.1 1108±45 1146±45 1236±66 2311±167 3056±45 1685±135
0.15 873±27 927±27 857±52 1710±64 2293±32 1083±49
Thyroid Nodules 0.05 3675±480 3577±413 4461±337 3513±212 6197±64 4068±388
0.1 2671±251 2645±216 3390±197 2762±129 3942±102 2826±200
0.15 2125±203 2127±203 2646±128 2355±131 2937±95 2188±167
Table 5: COMPASS produces the most efficient interval lengths on the majority of datasets across target coverages for SegResNet. For COMPASS-J, we use the final convolutional layer, which is the only layer in the segmentation head. Even without deeper layers, COMPASS still produces efficient intervals.

### v2

@@ C.3 Symmetric COMPASS
15:procedure Inference($x_{n+1}$, model, $h,\hat{\beta}$)
16:  Define $\hat{z}_{n+1}$ and compute $\Delta_{n+1}$ (as in Alg. 2).
@@
18:  return $S_{\hat{\beta}}(x_{n+1})$
19:end procedure
Algorithm 3 describes the symmetric formulation. During calibration, we compute a non-conformity score $R_{i}$ for each sample using a binary search (Algorithm 5). $R_{i}$ represents the minimal perturbation magnitude $\beta$ such that the interval $P(\hat{z}_{i},\Delta_{i},\beta)$ covers the true label $y_{i}$. We then compute the $(1-\alpha)$ quantile of these scores, $\hat{\beta}$. During inference, we apply this calibrated $\hat{\beta}$ to the test sample to produce the final prediction interval.
C.4 Asymmetric COMPASS
Algorithm 4 Asymmetric COMPASS (Calibration & Inference)
1:Calibration $D_{\mathrm{cal}}$, Test $x_{n+1}$, $\alpha_{\mathrm{lo}},\alpha_{\mathrm{hi}}$, Model $(f,g,V_{L})$.
@@ C.4 Asymmetric COMPASS
4:procedure Calibration($D_{\mathrm{cal}}$, model, $h,\alpha_{\mathrm{lo}},\alpha_{\mathrm{hi}}$)
5:  Initialize $\mathcal{R}_{\mathrm{lo}}\leftarrow\emptyset,\mathcal{R}_{\mathrm{hi}}\leftarrow\emptyset$.
6:  for $i=1,\dots,n$ do
7:   Define $\hat{z}_{i}$ and compute $\Delta_{i}$ (as in Alg. 2).
8:   $(R_{i,\mathrm{lo}},R_{i,\mathrm{hi}})\leftarrow\textsc{Asymmetric-Binary-Search}(\hat{z}_{i},\Delta_{i},y_{i})$.
9:   $\mathcal{R}_{\mathrm{lo}}\leftarrow\mathcal{R}_{\mathrm{lo}}\cup\{R_{i,\mathrm{lo}}\}$.
10:   $\mathcal{R}_{\mathrm{hi}}\leftarrow\mathcal{R}_{\mathrm{hi}}\cup\{R_{i,\mathrm{hi}}\}$.
11:  end for
12:  $\hat{\beta}_{\mathrm{lo}}\leftarrow\text{Quantile}(\mathcal{R}_{\mathrm{lo}},\frac{\lceil(1-\alpha_{\mathrm{lo}})(n+1)\rceil}{n})$.
13:  $\hat{\beta}_{\mathrm{hi}}\leftarrow\text{Quantile}(\mathcal{R}_{\mathrm{hi}},\frac{\lceil(1-\alpha_{\mathrm{hi}})(n+1)\rceil}{n})$.
14:  return $\hat{\beta}_{\mathrm{lo}},\hat{\beta}_{\mathrm{hi}}$
15:end procedure
16:
17:procedure Inference($x_{n+1}$, model, $h,\hat{\beta}_{\mathrm{lo}},\hat{\beta}_{\mathrm{hi}}$)
18:  Define $\hat{z}_{n+1}$ and compute $\Delta_{n+1}$ (as in Alg. 2).
19:  $S(x_{n+1})\leftarrow[h(g(\hat{z}_{n+1}-\hat{\beta}_{\mathrm{lo}}\Delta_{n+1})),h(g(\hat{z}_{n+1}+\hat{\beta}_{\mathrm{hi}}\Delta_{n+1}))]$.
20:  return $S(x_{n+1})$
21:end procedure
Algorithm 4 extends the framework to the asymmetric case. We independently calibrate a lower bound ($\hat{\beta}_{\mathrm{lo}}$) and an upper bound ($\hat{\beta}_{\mathrm{hi}}$) using separate non-conformity scores. This allows the interval to expand differently in the positive and negative directions, which is efficient for metrics with asymmetric sensitivity.
C.5 Score-Finding Procedures
@@
5:   $S_{\text{mid}}\leftarrow P(\hat{z},\Delta,b_{\text{mid}})$ $\triangleright$ Compute endpoint interval
6:   if $y\in S_{\text{mid}}$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
@@
10:  return $b_{\text{high}}$.
11:end function
Algorithm 6 Asymmetric Binary Search
1:function Asymmetric-Binary-Search($\hat{z},\Delta,y$)
2:  $m_{\text{lo}}(b)\leftarrow h(g(\hat{z}-b\Delta))$, $m_{\text{hi}}(b)\leftarrow h(g(\hat{z}+b\Delta))$
3:  Find Upper Bound Score ($R_{\mathrm{hi}}$):
4:  $b_{\text{low}}\leftarrow 0,b_{\text{high}}\leftarrow\beta_{\text{range}}$.
5:  for $k=1,\dots,k_{\text{max}}$ do
6:   $b_{\text{mid}}\leftarrow(b_{\text{low}}+b_{\text{high}})/2$.
7:   if $y\leq m_{\text{hi}}(b_{\text{mid}})$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
8:   else $b_{\text{low}}\leftarrow b_{\text{mid}}$
9:   end if
10:  end for
11:  $R_{\mathrm{hi}}\leftarrow b_{\text{high}}$.
12:  Find Lower Bound Score ($R_{\mathrm{lo}}$):
13:  $b_{\text{low}}\leftarrow 0,b_{\text{high}}\leftarrow\beta_{\text{range}}$.
14:  for $k=1,\dots,k_{\text{max}}$ do
15:   $b_{\text{mid}}\leftarrow(b_{\text{low}}+b_{\text{high}})/2$.
16:   if $y\geq m_{\text{lo}}(b_{\text{mid}})$ then $b_{\text{high}}\leftarrow b_{\text{mid}}$
17:   else $b_{\text{low}}\leftarrow b_{\text{mid}}$
18:   end if
19:  end for
20:  $R_{\mathrm{lo}}\leftarrow b_{\text{high}}$.
21:  return $(R_{\mathrm{lo}},R_{\mathrm{hi}})$.
22:end function
These are the root-finding functions necessary to determine the non-conformity scores. We use a binary search for computational efficiency. This requires $\mathcal{O}(\log N)$ forward passes, as opposed to $\mathcal{O}(N)$ for a linear search, making it highly scalable.
@@ Appendix D Additional Tables
Coverage (Mean±STD)
Dataset $\alpha$ COMPASS-J COMPASS-L E2E-CQR Local Output-CQR SCP
H&E 0.05 0.952±0.023 0.953±0.02 0.955±0.025 0.95±0.021 0.951±0.019 0.956±0.018
H&E 0.10 0.907±0.029 0.907±0.03 0.903±0.028 0.903±0.03 0.903±0.031 0.91±0.03
H&E 0.15 0.856±0.03 0.856±0.031 0.856±0.035 0.848±0.036 0.85±0.032 0.849±0.032
Skin Lesion 0.05 0.951±0.009 0.953±0.008 0.951±0.01 0.95±0.01 0.951±0.009 0.951±0.01
Skin Lesion 0.10 0.9±0.014 0.904±0.013 0.9±0.012 0.899±0.012 0.897±0.013 0.899±0.013
Skin Lesion 0.15 0.853±0.015 0.86±0.016 0.853±0.016 0.852±0.015 0.852±0.015 0.853±0.015
Nodule 0.05 0.949±0.018 0.953±0.018 0.958±0.014 0.948±0.017 0.954±0.016 0.953±0.019
Nodule 0.10 0.901±0.025 0.904±0.026 0.904±0.022 0.906±0.025 0.908±0.024 0.906±0.024
Nodule 0.15 0.852±0.032 0.858±0.031 0.849±0.03 0.85±0.028 0.845±0.026 0.857±0.027
PolyP 0.05 0.955±0.02 0.96±0.018 0.957±0.019 0.951±0.022 0.955±0.018 0.96±0.022
PolyP 0.10 0.899±0.034 0.908±0.032 0.902±0.027 0.897±0.032 0.901±0.029 0.907±0.03
PolyP 0.15 0.849±0.035 0.865±0.03 0.853±0.034 0.854±0.037 0.856±0.033 0.856±0.032
Interval Size (Pixels2, Mean±STD)
@@
H&E 0.05 4637±630 4408±432 9702±1535
H&E 0.10 3160±336 3139±375 6463±1193
H&E 0.15 2320±252 2354±146 4979±647
Skin Lesion 0.05 1657±80 1689±83 14838±10
Skin Lesion 0.10 1179±53 1208±58 12445±51
Skin Lesion 0.15 934±30 956±33 8267±79
Nodule 0.05 3257±210 3394±280 16247±2
Nodule 0.10 2444±174 2510±180 16143±3
Nodule 0.15 2016±143 2082±142 16056±4
PolyP 0.05 5489±575 6376±769 15937±9
PolyP 0.10 4056±293 4397±469 15644±10
PolyP 0.15 3394±290 3686±361 15353±14
Table 4: Comparison of layer choice. For $\alpha=0.1$, we find that while for some datasets the efficiency gains are comparable, penultimate features generally achieve shorter interval lengths.
Interval Size (Pixels2,Mean±STD)
Dataset COMPASS-J (Bottleneck) COMPASS-J (Deep) COMPASS-J (Shallow) COMPASS-L
H&E 9394±799 3160±336 3140±386 3139±375
Skin Lesion 4084±154 1179±53 1210±51 1208±58
Nodule 8070±494 2444±174 2500±187 2510±180
PolyP Non-monotonic 4056±293 4222±390 4397±469
@@
H&E Class COMPASS-J 2055±220 0.911±0.024
COMPASS-L 2065±263 0.903±0.024
E2E-CQR 2444±278 0.900±0.027
Local 3113±201 0.907±0.024
Output-CQR 2890±265 0.901±0.024
SCP 2371±290 0.900±0.026
Feature COMPASS-J 1916±177 0.909±0.024
COMPASS-L 1944±244 0.896±0.024
E2E-CQR 2197±288 0.884±0.029
Jacobian COMPASS-J 1822±200 0.896±0.031
Unweighted COMPASS-J 1810±127 0.886±0.025
COMPASS-L 1759±159 0.884±0.024
E2E-CQR 1993±196 0.871±0.026
Local 2833±158 0.88±0.027
Output-CQR 2563±134 0.879±0.024
SCP 2003±169 0.878±0.024
Skin Lesion Class COMPASS-J 1318±132 0.896±0.019
COMPASS-L 1373±136 0.895±0.019
E2E-CQR 1335±277 0.887±0.021
Local 2426±354 0.893±0.02
Output-CQR 3380±43 0.888±0.018
SCP 1905±475 0.888±0.024
Feature COMPASS-J 1573±158 0.919±0.014
COMPASS-L 1669±168 0.919±0.013
E2E-CQR 2030±269 0.919±0.012
Jacobian COMPASS-J 1485±126 0.914±0.013
Unweighted COMPASS-J 2183±132 0.94±0.005
COMPASS-L 2211±150 0.939±0.007
E2E-CQR 3400±226 0.945±0.006
Local 5446±380 0.947±0.006
Output-CQR 4201±145 0.945±0.008
SCP 4707±323 0.948±0.006
Table 6: COMPASS produces the most efficient interval lengths on the majority of datasets across target coverages for SegResNet.
Interval Size (Pixels2, Mean±STD)